A random sample of size 20 is selected from a normal population, where the population standard deviation is unknown. Assume that a two-tailed test at the 1% significance level is to be used. For what value of t will the null hypothesis not be rejected?

Select one:

a. None of the suggested answers are correct.

b. to the left of -2.845 or to the right of 2.845

c. to the left of -2.528 or to the right of 2.528

d. to the left of -2.539 or to the right of 2.539

e. to the left of -2.861 or to the right of 2.861
Answer #1

Solution :

Given that ,

n = 20

df = 20 - 1 = 19

\alpha / 2 = 0.005

t 0.005,19 = +/-2.861

e. to the left of -2.861 or to the right of 2.861